OpenStudy (anonymous):

Find the vertex of the parabola whose equation is y = x^2 - 4x + 6. Find the y-coordinate of the vertex of the parabola whose equation is y = x^2 - x + 2.

OpenStudy (anonymous):

and once again, \[-\frac{b}{2a}\]

OpenStudy (anonymous):

\[ y = x^2 - 4x + 6\] \[a=1,b=-4\]

OpenStudy (anonymous):

let me know what you get

OpenStudy (sweetburger):

you could always go full derp and use a graphing calculator but it is far easier to just use the formula

OpenStudy (anonymous):

4/2 which is 2?

OpenStudy (anonymous):

yes, right that is the first coordinate of the vertex

OpenStudy (anonymous):

whats the equation for the second coordinate?

OpenStudy (anonymous):

second coordinate is \[ y = 2^2 - 4\times 2 + 6\]

OpenStudy (anonymous):

i.e. what you get for \(y\) when you replace \(x\) by the first coordinate

OpenStudy (anonymous):

in this case it is \(y=4-8+6=-4+6=2\) so both the first and second coordinate of the vertex is \(2\) and the vertex is \((2,2)\)

OpenStudy (anonymous):

is the y coordinate the first or second one in the parentheses?

OpenStudy (anonymous):

Find the y-coordinate of the vertex of the parabola whose equation is \(y = x^2 - x + 2\) is a bit trickier because you have to work with fractions

OpenStudy (anonymous):

the \(y\) coordinate is the second one it goes \((x,y)\)

OpenStudy (anonymous):

easy enough to remember because \(x\) comes before \(y\) in the alphabet

OpenStudy (anonymous):

ah, then i think the answer is 1 and 3/4

OpenStudy (anonymous):

for \[y=x^2-x+2\]?

OpenStudy (anonymous):

oh wait wait wait, is it 2 and 1/4?

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

2? -.-

OpenStudy (anonymous):

it is still \[-\frac{b}{2a}\] but in this case \(a=1,b=-1\) so the first coordinate of the vertex is \(\frac{1}{2}\) not \(2\) you had it upside down

OpenStudy (anonymous):

the second coordinate takes some arithmetic it is \[y=\left(\frac{1}{2}\right)^2-\frac{1}{2}+2\]

OpenStudy (anonymous):

which, if my arithmetic is correct, is \[\frac{1}{4}-\frac{1}{2}+2=-\frac{1}{4}+2=\frac{7}{4}\]

OpenStudy (anonymous):

wanna check it?

OpenStudy (anonymous):

which is 1 and 3/4!

OpenStudy (anonymous):

i knew that was the right answer!

OpenStudy (anonymous):

hold the phone it is \((\frac{1}{2},\frac{7}{4})\) which is not \((1,\frac{3}{4})\)

OpenStudy (anonymous):

clear?

OpenStudy (anonymous):

no, 7/4 = 1 and 3/4

OpenStudy (anonymous):

i only need the y

OpenStudy (anonymous):

right dad?

OpenStudy (anonymous):

ooh i see what you are saying!!

OpenStudy (anonymous):

yes dear now don't stay out too late

OpenStudy (anonymous):

lol! i wont, im just going to get pellet faced and be a whore!

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

you done? or are there more

OpenStudy (anonymous):

ill make a new post, and omg thanks for helping me so much dad!

OpenStudy (anonymous):

yw dear